OpenStudy (anonymous):

A triangular prism has a surface area of 264 square feet, a length of 15 feet, a height of 4 feet, and sides of 5 feet. Find the length of the base (width) of the triangular prism. please help i'm homeschooled and I need to finish this test. Thank you!

OpenStudy (jadedry):

If its sides are 5 feet, then it must be have isosceles sides. Something like this: http://everythingmaths.co.za/maths/grade-10/12-measurement/pspictures/615bfe0f93eff216de1859c123a71176.png To find the width, you'll have to use Pythagoras' theorem. so: \[\sqrt{5^2 - 4^2}\] = width * 1/2 = 3 therefore width = 6.

OpenStudy (anonymous):

Thank you so much!!!

OpenStudy (jadedry):

No problem, good luck with your studies! c: